OpenStudy (anonymous):

I need some help. I am having a hard time understanding and solving radical expressions. Can anyone give me a step by step explanation of how to solve them?

OpenStudy (anonymous):

It would help if you write a problem. It is hard to talk about these things abstractly.

OpenStudy (anonymous):

This is what we are working on and I am completely lost; \[\sqrt5 + \sqrt18\]

OpenStudy (anonymous):

There is not much more that can be done with this. The 'base' must be the same. The 5 and 18 are different.

OpenStudy (anonymous):

hmmm...I am confused then! :) because this is what my classmates are posting and they are solving them

OpenStudy (anonymous):

This is one for example: √30-√49

OpenStudy (anonymous):

\[\sqrt{30}-7\]

OpenStudy (anonymous):

Is that the final answer?

OpenStudy (anonymous):

yeah, your first example is really\[5^{1/2}+18^{1/2}\]It is like\[a ^{1/2}+b ^{1/2}\]You can't really add the exponents of a and b because the bases are different.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

My point stands, but in math one has to be careful and precise with language. You don't 'add' the exponents unless it is a multiplication operation, eg,\[a ^{1/2}a ^{1/2}=a ^{1} \]\[a ^{1/2}+a ^{1/2}=2a ^{1/2}\]